OpenStudy (anonymous):

(2x+3)+(2y-2)y'=0 differential problem

OpenStudy (ash2326):

we have \[2x+3+(2y-2)\frac{dy}{dx}=0\]we can write this as \[(2x+3)dx=-(2y-2)dy\] let's integrate both sides \[x^2+3x+C= -(y^2-2y)\] so the differential equation's solution is \[y^2-2y+x^2+3x+C=0\] C is constant of integration

OpenStudy (ash2326):

Did you understand Gia?

OpenStudy (anonymous):

yes, thanks

OpenStudy (ash2326):

welcome:)

OpenStudy (anonymous):

how about this (2xy^2+2y) + (2x^2 y + 2x) y' = 0

OpenStudy (ash2326):

Gia I'm gonna solve this but next time use one post for only one question

OpenStudy (anonymous):

okay, i'm newbee in here, sorry if i did a mistake

OpenStudy (ash2326):

No, problem Gia :D Welcome to Open Study

OpenStudy (ash2326):

We have \[(2xy^2+2y) + (2x^2 y + 2x) \frac{dy}{dx} = 0 \] writing this as \[(2xy^2+2y)dx+(2x^2y+2x)dy\] If you notice the terms the first term is the derivative of \(x^2y^2+2xy\) with respect to x and the second term is the derivative of \(x^2y^2+2xy\) with respect to y. So this can be written as \[d(x^2y^2+2xy)=0\] you do know that \[ d(xy)= ydx+xdy\] and \[d(x^2y^2)=2xy^2dx+2x^2ydy\] therefore we have \[d(x^2y^2+2xy)=0\] integrate both sides, you'll get \[x^2y^2+2xy=C\]